# Efficient first-principles inverse design of nanolasers

### v2

@@ 1 Introduction
[…]
@@
A key enabling factor for first-principles nanolaser design is the recent development of SALT, reviewed in Sec. 2.1 [1, 2], which formulates a tractable model of a full nonlinear lasing problem by transforming the time-domain nonlinear Maxwell–Bloch equations [34, 35] into a set of self-consistent, time-independent equations for the lasing steady state. SALT has provided a foundation for many subsequent developments, such as improvements in the theory of laser amplifiers [36], complicated gain media with carrier diffusion [4], quantum limits on the laser linewidth [37], and random lasers [38]. However, the fact that SALT is based on a nonlinear eigenproblem (arising when a resonance lases) makes it more challenging for inverse design than linear scattering: not only does it have all of the pole-tracking and potential non-differentiability of linear eigenvalue/resonance optimization [21], but when a pole crosses the real-$\omega$ axis one must also “turn on” a nonlinear solver [39]. Merely identifying the lasing pole in a differentiable way would be difficult without a good starting guess (perhaps via LDOS optimization), and the nonlinear effects incur additional computational complexity [39]. Although these issues may be tractable in principle (e.g. differentiating the nonlinear solve via implicit adjoint methods [40]; such methods have been used to optimize nonlinear steady-state transmission [41]), it would be better to find a more efficient formulation. The reason such an improvement should be possible is the fact that an optimized nanolaser is quite special: to be high-efficiency and/or low threshold, the cavity will have a high $Q$, and it is known that SALT can be simplified by perturbative analysis in this regime [2] (empirically found to be accurate for passive cavity resonances with $Q\gtrsim 100$ [42]), although the previous perturbative “SPA-SALT” techniques were not devised with optimization in mind. Here, we show that specifically targeting efficient inverse design allows us to obtain a computationally efficient and simple FOM without sacrificing first-principles physics.
@@ 2.1 Review of Steady-State Ab-initio Laser Theory (SALT)
In a laser, an external pump excites a gain medium, which then interacts with the optical cavity mode, and emits photons into an output mode through stimulated emission. This is illustrated in Fig. 1(a) for an example of an inverse designed nanolaser. The gain medium located in the cavity is modeled as an ensemble of two-level atoms interacting with the optical cavity field, which can be described by means of the Maxwell-Bloch equations [34, 35, 1]
××$\begin{gathered}\nabla\times\nabla\times\mathbf{E}^{+}-\frac{\varepsilon_{c}}{c^{2}}\ddot{\mathbf{E}}^{+}=\frac{1}{c^{2}\varepsilon_{0}}\ddot{\mathbf{P}}^{+}\,,\\ \dot{\mathbf{P}}^{+}=\left(-i\omega_{a}-\gamma_{\perp}\right)\mathbf{P}^{+}+\frac{g^{2}}{i\hbar}\mathbf{E}^{+}D\,,\\ \dot{D}=\gamma_{\|}\left(D_{0}-D\right)-\frac{2}{i\hbar}\left[\mathbf{E}^{+}\cdot\left(\mathbf{P}^{+}\right)^{*}-\mathbf{P}^{+}\cdot\left(\mathbf{E}^{+}\right)^{*}\right]\,,\end{gathered}$ (1)
where $i$ is the imaginary unit, $\hbar$ is the reduced Planck constant, $g$ is a dipole matrix element, $\mathbf{E}^{+}(\mathbf{r},t)$ and $\mathbf{P}^{+}(\mathbf{r},t)$ are the positive frequency components of the electric field and polarizations respectively, $\varepsilon_{c}(\mathbf{r})$ is the complex relative permittivity of the cavity, $\varepsilon_{0}$ is the vacuum permittivity, $D(\mathbf{r},t)$ is the population inversion, $D_{0}(\mathbf{r})$ is the gain profile, which is established through external pumping, $\omega_{a}$ is the atomic transition frequency, $\gamma_{\perp}$ is the gain width (polarization dephasing rate), and $\gamma_{\parallel}$ is the population relaxation rate. Note that these equations rely on the rotating-wave approximation (RWA) [2].
For semiconductor lasers, the population relaxation rate will be much lower than the atomic transition frequency and the dephasing rate ($\omega_{a},\gamma_{\perp}\gg\gamma_{\parallel}$), which leads to the stationary-inversion approximation (SIA) $\dot{D}\sim 0$ [2]. One can then derive the coupled nonlinear SALT differential equations [2]
[(××${\left[\left(\nabla\times\nabla\times{}\right)-\left(\varepsilon_{c}(\mathbf{r})+\frac{\gamma_{\perp}D(\mathbf{r})}{\omega_{i}-\omega_{a}+i\gamma_{\perp}}\right)\left(\frac{\omega_{i}}{c}\right)^{2}\right]\mathbf{E}_{i}(\mathbf{r})=0}\,,$ (2)
where $\omega_{i}$ and $\mathbf{E}_{i}$ are the frequency and mode profile of the $i$-th cavity mode, and the expression for the population inversion is given by
@@
where $d$ is a scalar-valued pumping strength that we have separated from the spatial pumping profile $D_{0}(\mathbf{r})$, $\Gamma_{j}=\gamma_{\perp}^{2}/\left[\gamma_{\perp}^{2}+(\omega_{j}-\omega_{a})^{2}\right]$ is a Lorentzian gain curve evaluated at the frequency $\omega_{j}$ for a gain medium emitting at the wavenumber $\omega_{a}$, and $e_{c}$ is a non-dimensionalization scaling for the fields [2].
@@ 2.2 Single-pole threshold approximation
The expression in Eq. (2) simplifies when considering a single high-$Q$ lasing mode $L$ ($i=j=L$) that is aligned to the atomic transition $\omega_{a}=\omega_{L}$ of the emitters in the gain medium. This approximation, which is accurate for a sufficiently high $Q$ (e.g. $Q\gtrsim 100$ in previous work [42]) is also known as the single-pole approximation (SPA-) SALT equation [2]
[(××${\left[(\nabla\times\nabla\times)-\left(\varepsilon_{c}(\mathbf{r})-i\Delta\varepsilon_{\Im}(\mathbf{r})\right)\left(\frac{\omega_{L}}{c}\right)^{2}\right]\mathbf{E}_{L}(\mathbf{r})=0}\,.$ (4)
@@
Eq. (4) still involves solving one nonlinear eigenvalue problem. Nevertheless, the expression can be further simplified by perturbatively reducing the equation to linear form. For a high-$Q$ cavity, operating at low powers near threshold, the gain required to balance cavity loss is small. More precisely, in this regime the change in the imaginary part of the dielectric permittivity induced by the active medium will be small ($\propto 1/Q$) with respect to the passive cavity permittivity ($\Delta\varepsilon_{\Im}\ll\varepsilon_{c}$), for a high-$Q$ passive cavity mode with a complex angular frequency frequency $\omega_{\mathrm{m}}=\omega_{0}-i\gamma_{0}$, where the quality factor is defined as $Q=\omega_{0}/(2\gamma_{0})$. We can then describe the shift in the cavity-mode frequency by using first-order perturbation theory [43]
$\Delta\omega_{\mathrm{m}}=i\frac{\omega_{0}}{2}\frac{\int_{\Omega}\Delta\varepsilon_{\Im}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,,$ (6)
where $\mathbf{E}_{\mathrm{m}}$ is the mode profile of the passive cavity. In the high-$Q$ cavity limit, the lasing mode $\mathbf{E}_{L}$ is approximately proportional to the cavity mode [2] within a $\mathcal{O}(1/\sqrt{Q})$ error (proportional to the square root of the radiated power [3]). The change in permittivity can then be expressed as
$\Delta\varepsilon_{\Im}(\mathbf{r})\approx\frac{D_{0}(\mathbf{r})d}{1+|a_{0}|^{2}\,\left(e_{c}^{-2}|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}/\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega\right)}\,,$ (7)
where $a_{0}$ is the amplitude of the lasing mode and the lasing intensity $|a_{0}|^{2}$ has units of energy. Note that since $\Delta\varepsilon_{\Im}\in\mathbb{R}$ this yields a purely imaginary shift in the cavity-mode frequency, which can be used to counteract the losses of the passive cavity, so that the cavity mode starts lasing.
In the low-power regime, for a small amplitude $|a_{0}|^{2}$ of the lasing mode, one can expand Eq. (7) in a Taylor series around $|a_{0}|^{2}=0$. As shown in Fig. 1(b) the amplitude of the lasing mode is zero at threshold, and the system will lase for a pumping strength above threshold $d\geq d_{\text{thresh}}$. Notice, that we are here considering the approximation where we neglect the spontaneous emission into the lasing mode, which will lead to non-zero power in the cavity mode below the lasing threshold [10]. In the zero-amplitude limit, the change in permittivity is linearly related to the pump strength: $\Delta\varepsilon_{\Im}(\mathbf{r})=d\,D_{0}(\mathbf{r})$. As shown in Fig. 1(c), for lasing to occur, the imaginary frequency shift caused by the gain, as expressed in Eq. (6), must balance the losses of the passive cavity
$\Im\{\Delta\omega_{\mathrm{m}}\}-\gamma_{0}=0\,.$ (8)
@@
$d_{\text{thresh}}=\frac{1}{\Gamma Q}=\frac{1}{Q}\frac{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,.$ (9)
As would be expected, the pumping strength needed to reach the lasing threshold is inversely proportional to $Q$ and to the energy confinement in the active medium $\Gamma$, which leads to stronger light-matter interaction. In App. A, we show how the same result can be recovered using SPA-SALT. Note that in the single-emitter limit where $D_{0}(\mathbf{r})\propto\delta(\mathbf{r}-\mathbf{r}^{\prime})$, Eq. (9) simplifies to $d_{\text{thresh}}=V/Q$, where $V$ is the typical “mode volume” of the resonance, recovering an expression inversely proportional to the Purcell enhancement and the LDOS [6, 7, 8]; we comment further on this regime in Sec. 2.3.4.
Once the pumping strength crosses the lasing threshold, the cavity-mode will start to lase with an amplitude $a_{0}$, as depicted in Fig. 1(b). Considering both the zeroth-order and the first-order correction in the small amplitude expansion of Eq. (7), we derive an approximate expression for the lasing intensity
$|a_{0}|^{2}\approx\frac{\Delta d}{d_{\text{thresh}}}\frac{{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{e_{c}^{-2}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{E}_{\mathrm{m}}(\mathbf{r})|^{4}\,\mathop{}\!\mathrm{d}\Omega}\,,$ (10)
where we have assumed that we are operating in the weak pumping limit (just above threshold), where $d\approx d_{\text{thresh}}$ and where $\Delta d=d-d_{\text{thresh}}\ll 1$ represents the pumping strength above threshold. For more details on this derivation, consult App. B. Note that evaluating Eqs. (910) still requires one to solve a linear eigenproblem to determine the modes of the passive cavity, as in SPA-SALT. Nevertheless, assuming a high-$Q$ cavity $(Q\gtrsim 100$), we can eliminate this requirement by exciting the field in the cavity with an external source (e.g., a waveguide coupled to the cavity), since the field in the cavity will be approximately proportional to the passive cavity mode $\mathbf{E}_{\mathrm{m}}$ [3, 44]. Although this eliminates the requirement of an eigensolve, it does not include information about the efficiency of coupling between the cavity mode and the output channel. In the next section, we will use TCMT to encode the out-coupling into a single, convenient FOM.

### v3

@@ 1 Introduction
Figure 1: (a) Schematic nanolaser: pump power $P_{\text{pump}}$ excites a gain profile $D_{0}$ and a lasing mode that outputs power $P_{\text{out}}$ into a channel (waveguide), for a cavity (from Sec. 3.1) that maximizes efficiency $\sim P_{\text{out}}/P_{\text{pump}}$. (b) Lasing amplitude $|a_{0}|^{2}$ ($\sim\text{energy}$) vs. pump strength $d$, lasing above a threshold $d_{\text{thresh}}$. (c) Change $\Im\{\Delta\omega\}$ in modal gain/loss depends nonlinearly on amplitude $|a|^{2}$, and stabilizes at $|a_{0}|^{2}$, where $\Im\{\Delta\omega\}=$ passive cavity loss $\gamma_{0}$.
In this paper, we present a new formulation for efficient inverse design of nanolasers (Fig. 1) to maximize efficiency ($\sim\text{output}$ $\text{power}/\text{pump}$ power), taking nonlinear and coupling effects into account, and demonstrate 2d and 3d semiconductor-laser topology optimization (TopOpt, Sec. 2.5). We exploit the fact that optimization evolves the lasing cavity to a high-$Q$ ($\gtrsim 100$, long-lifetime) regime to perturbatively extract hole-burning effects, thresholds, output coupling, and other relevant phenomena from a single linear “reciprocal” scattering solve excited from the output port for a passive cavity. Our approach builds on the “SALT” (steady-state ab initio laser theory) formulation of lasing [SALT_original, Ge_2010] (Sec. 2.1), which exactly solves the nonlinear Maxwell–Bloch equations for a single-mode lasing steady state, including all nonlinear gain and saturation effects in the Maxwell–Bloch model, but we are able to eliminate expensive nonlinear and linear eigensolves using a new synthesis of SALT, perturbation theory (extending SPA-SALT theory [Ge_2010], Sec. 2.2), and temporal coupled-mode theory (TCMT [phot_crys], Sec. 2.3). Our resulting figure of merit (FOM) is a simple ratio $\sim(\int_{\text{gain}}|\mathbf{E}|^{2}\mathop{}\!\mathrm{d}\Omega)^{3}/\int_{\text{gain}}|\mathbf{E}|^{4}\mathop{}\!\mathrm{d}\Omega$ (Eq. (14) in Sec. 2.3.3) computed from only the linear-model electric field $\mathbf{E}$. We show that this FOM reduces to a more conventional cavity LDOS-like optimization—power emitted into an output channel by a dipole—in the limit of a point-like gain region (Sec. 2.3.4), but leads to very different (more delocalized) cavity designs and performance for a distributed gain region. We demonstrate TopOpt of this FOM with both 2d and 3d semiconductor-on-insulator designs, including manufacturing constraints, and investigate the effect of gain-region diameter (Sec. 3.1). We also compare to a heuristic FOM $\sim\int_{\text{gain}}|\mathbf{E}|^{2}\mathop{}\!\mathrm{d}\Omega$ (Sec. 2.3.4), and show that this leads to suboptimal performance (Sec. 3.1). Furthermore, we show how to incorporate gain-diffusion effects [csalt] into TopOpt (e.g., for semiconductor lasers in a free-carrier approximation) at the cost of only two additional damped-diffusion solves (Sec. 2.4), and show that diffusion favors a disconnected topology of the gain region (Sec. 3.2). We believe that our results not only demonstrate the importance (and surprising ease) of incorporating sophisticated nonlinear lasing models into the inverse-design process, but also provide a foundation for future extensions, e.g. to include an explicit model of the pumping process, as discussed in Sec. 4.
The inverse design of nanolasers has conventionally been based on improving generic FOMs that are related to lasing, without directly modeling the laser physics or accounting for the spatial distribution of the emitters [xiong2025nanolaser], such as maximizing the cavity quality factor $Q$ ($\sim\text{lifetime}/\text{period}$ [phot_crys]). Another such FOM is the local density of states (LDOS), which is proportional to $Q$ divided by a measure of modal volume (the “Purcell factor”) and is more precisely the total power expended by a dipole source [OskooiJo13-sources]; this is related to enhancement of light–matter interactions (e.g. spontaneous emission rate) at a single point [LDOS_Laser, LDOS_Laser_2, LDOS_Laser_3, bowtie_laser]. For example, high $Q$ factors lead to lower lasing thresholds [saldutti2023thresholdsemiconductornanolasers] (Sec. 2.2). Nonetheless, and as our results also demonstrate, in nanolaser design a high $Q$ alone is insufficient: while strong temporal confinement (high $Q$) is crucial, the photons emitted by the gain medium must also be efficiently extracted and coupled into an output channel. This concept has been explored in the context of inverse design of light emission, where the optical extraction efficiency [extraction, Janssen2010, Mayer, Chakravarthi:20, Yao_2022, Chung:22, KimShinLeeParkLeeParkSeoParkParkJang, Shultzman23, schubert2023fouriermodalmethodinverse] —the (linear-model) output power of some source into some channel(s)—is considered as the optimization FOM. However, for laser design, simple metrics such as the $Q$, LDOS, or extraction efficiency do not account for the precise distribution of the gain medium and its overlap with the optical field [xiong2025nanolaser] or nonlinear lasing effects (e.g. gain, saturation). Below, we show that an LDOS-like FOM only becomes equivalent to a more complete lasing model in the limiting case of a point-like gain region (as shown in Sec. 2.3.4). The LDOS is also relevant to other physical phenomena such as sensing, and there have been several research efforts targeting this FOM using inverse-design frameworks like TopOpt [LDOS_opt_wang, LDOS_opt_liang, LDOS_bounds, albrechtsen_nanometer-scale_2022, LDOS_G], including experimental realization [albrechtsen_nanometer-scale_2022], along with other methods targeting high QQ[Ahn2022]. (TopOpt, reviewed in Sec. 2.5, is a gradient-based inverse design framework that finds freeform designs via large-scale optimization where “every pixel” is a degree of freedom, and has been successfully applied to many photonic problems [jensen_topology_2011, Molesky_2018].) However, it is known that the result of cavity optimization can be greatly altered if one targets an FOM that is tailored to a phenomenon very different from spontaneous emission at a point, such as Raman scattering [ChristiansenMi20, PanCh21, YaoVe23], volume-averaged emission [Yao_2022], scintillation effects [Roques_Carmes_2022], or harmonic generation [LinLi16, Mann2023].
@@
A key enabling factor for first-principles nanolaser design is the recent development of SALT, reviewed in Sec. 2.1 [SALT_original, Ge_2010], which formulates a tractable model of a full nonlinear lasing problem by transforming the time-domain nonlinear Maxwell–Bloch equations [haken_Laser_dynamics, PhysRev.134.A1429] into a set of self-consistent, time-independent equations for the lasing steady state. SALT has provided a foundation for many subsequent developments, such as improvements in the theory of laser amplifiers [I-SALT], complicated gain media with carrier diffusion [csalt], quantum limits on the laser linewidth [pick], and random lasers [random_lasers]. However, the fact that SALT is based on a nonlinear eigenproblem (arising when a resonance lases) makes it more challenging for inverse design than linear scattering: not only does it have all of the pole-tracking and potential non-differentiability of linear eigenvalue/resonance optimization [LDOS_opt_liang], but when a pole crosses the real-$\omega$ axis one must also “turn on” a nonlinear solver [EsterhazyLiu14]. Merely identifying the lasing pole in a differentiable way would be difficult without a good starting guess (perhaps via LDOS optimization), and the nonlinear effects incur additional computational complexity [EsterhazyLiu14]. Although these issues may be tractable in principle (e.g. differentiating the nonlinear solve via implicit adjoint methods [Griewank2008]; such methods have been used to optimize nonlinear steady-state transmission [Hughes2018]), it would be better to find a more computationally efficient formulation. The reason such an improvement should be possible is the fact that an optimized nanolaser is quite special: to be high-efficiency and/or low threshold, the cavity will have a high $Q$, and it is known that SALT can be simplified by perturbative analysis in this regime [Ge_2010] (empirically found to be accurate for passive cavity resonances with $Q\gtrsim 100$ [cerjan_2016] and numerically validated in App. D), although the previous perturbative “SPA-SALT” techniques were not devised with optimization in mind. Here, we show that specifically targeting efficient inverse design allows us to obtain a computationally efficient and simple FOM without sacrificing first-principles physics.
@@ 2.1 Review of Steady-State Ab-initio Laser Theory (SALT)
In a laser, an external pump excites a gain medium, which then interacts with the optical cavity mode, and emits photons into an output mode through stimulated emission. This is illustrated in Fig. 1(a) for an example of an inverse designed nanolaser. The gain medium located in the cavity is modeled as an ensemble of two-level atoms interacting with the optical cavity field, which can be described by means of the Maxwell-Bloch equations [haken_Laser_dynamics, PhysRev.134.A1429, SALT_original]
[…]
where $i$ is the imaginary unit, $\hbar$ is the reduced Planck constant, $g$ is a dipole matrix element, $\mathbf{E}^{+}(\mathbf{r},t)$ and $\mathbf{P}^{+}(\mathbf{r},t)$ are the positive frequency components of the electric field and polarizations respectively, $\varepsilon_{c}(\mathbf{r})$ is the complex relative permittivity of the cavity, $\varepsilon_{0}$ is the vacuum permittivity, $D(\mathbf{r},t)$ is the population inversion, $D_{0}(\mathbf{r})$ is the gain profile, which is established through external pumping, $\omega_{a}$ is the atomic transition frequency, $\gamma_{\perp}$ is the gain width (polarization dephasing rate), and $\gamma_{\parallel}$ is the population relaxation rate. Note that these equations rely on the rotating-wave approximation (RWA) [Ge_2010].
For semiconductor lasers, the population relaxation rate will be much lower than the atomic transition frequency and the dephasing rate ($\omega_{a},\gamma_{\perp}\gg\gamma_{\parallel}$), which leads to the stationary-inversion approximation (SIA) $\dot{D}\sim 0$ [Ge_2010]. One can then derive the coupled nonlinear SALT differential equations [Ge_2010]
$\left[\left(\nabla\times\nabla\times{}\right)-\left(\varepsilon_{c}(\mathbf{r})+\frac{\gamma_{\perp}D(\mathbf{r})}{\tilde{\omega}_{i}-\omega_{a}+i\gamma_{\perp}}\right)\left(\frac{\tilde{\omega}_{i}}{c}\right)^{2}\right]\mathbf{E}_{i}(\mathbf{r})=0$}\,, (2)
where $\omega_{i}$ and $\mathbf{E}_{i}$ are the complex-valued frequency and mode profile of the $i$-th cavity mode, and the expression for the population inversion is given by
@@
where $d$ is a scalar-valued pumping strength that we have separated from the spatial pumping profile $D_{0}(\mathbf{r})$, $\Gamma_{j}=\gamma_{\perp}^{2}/\left[\gamma_{\perp}^{2}+(\omega_{j}-\omega_{a})^{2}\right]$ is a Lorentzian gain curve evaluated at the frequency $\omega_{j}$ for a gain medium emitting at the wavenumber $\omega_{a}$, and $e_{c}$ is a non-dimensionalization scaling for the fields [Ge_2010].
@@ 2.2 Single-pole threshold approximation
The expression in Eq. (2) simplifies when considering a single high-$Q$ lasing mode $L$ ($i=j=L$) that is aligned to the atomic transition $\omega_{a}=\omega_{L}$ of the emitters in the gain medium. This approximation, which is accurate for a sufficiently high $Q$ (e.g. $Q\gtrsim 100$ in previous work [cerjan_2016] and numerical validation in App. D) is also known as the single-pole approximation (SPA-) SALT equation [Ge_2010]
$\left[(\nabla\times\nabla\times)-\left(\varepsilon_{c}(\mathbf{r})-i\Delta\varepsilon_{\Im}(\mathbf{r})\right)\left(\frac{\omega_{L}}{c}\right)^{2}\right]\mathbf{E}_{L}(\mathbf{r})=0$}\,. (4)
@@
Eq. (4) still involves solving one nonlinear eigenvalue problem. Nevertheless, the expression can be further simplified by perturbatively reducing the equation to linear form. For a high-$Q$ cavity, operating at low powers near threshold, the gain required to balance cavity loss is small. More precisely, in this regime the change in the imaginary part of the dielectric permittivity induced by the active medium will be small ($\propto 1/Q$) with respect to the passive cavity permittivity ($\Delta\varepsilon_{\Im}\ll\varepsilon_{c}$), for a high-$Q$ passive cavity mode with a complex angular frequency frequency $\tilde{\omega}_{1}=\omega_{1}-i\gamma_{1}$, where the quality factor is defined as $Q=\omega_{1}/(2\gamma_{1})$. We can then describe the shift in the cavity-mode frequency by using first-order perturbation theory [Raman_2011]
$\Delta\tilde{\omega}_{1}=i\frac{\omega_{1}}{2}\frac{\int_{\Omega}\Delta\varepsilon_{\Im}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,$, (6)
where $\mathbf{e}_{1}$ is the mode profile of the passive cavity. In the high-$Q$ cavity limit, the lasing mode $\mathbf{E}_{L}$ is approximately proportional to the cavity mode [Ge_2010] within a $\mathcal{O}(1/\sqrt{Q})$ error (proportional to the square root of the radiated power [phot_crys]). The change in permittivity can then be expressed as
$\Delta\varepsilon_{\Im}(\mathbf{r})\approx\frac{D_{0}(\mathbf{r})d}{1+|a_{1}|^{2}\,\left(e_{c}^{-2}|\mathbf{e}_{1}(\mathbf{r})|^{2}/\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega\right)}\,$, (7)
where $a_{1}$ is the amplitude of the lasing mode and the lasing intensity $|a_{1}|^{2}$ has units of energy. Note that since $\Delta\varepsilon_{\Im}\in\mathbb{R}$ this yields a purely imaginary shift in the cavity-mode frequency, which can be used to counteract the losses of the passive cavity, so that the cavity mode starts lasing.
In the low-power regime, for a small amplitude $|a_{1}|^{2}$ of the lasing mode, one can expand Eq. (7) in a Taylor series around $|a_{1}|^{2}=0$. As shown in Fig. 1(b) the amplitude of the lasing mode is zero at threshold, and the system will lase for a pumping strength above threshold $d\geq d_{\text{thresh}}$. Notice, that we are here considering the approximation where we neglect the spontaneous emission into the lasing mode, which will lead to non-zero power in the cavity mode below the lasing threshold [saldutti2023thresholdsemiconductornanolasers]. In the zero-amplitude limit, the change in permittivity is linearly related to the pump strength: $\Delta\varepsilon_{\Im}(\mathbf{r})=d\,D_{0}(\mathbf{r})$. As shown in Fig. 1(c), for lasing to occur, the imaginary frequency shift caused by the gain, as expressed in Eq. (6), must balance the losses of the passive cavity
$\Im\{\Delta\tilde{\omega}_{1}\}-\gamma_{1}=0\,.$ (8)
@@
$d_{\text{thresh}}=\frac{1}{Q}\frac{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}\,.$ (9)
As would be expected, the pumping strength needed to reach the lasing threshold is inversely proportional to $Q$ and to the energy confinement in the active medium, which leads to stronger light-matter interaction. In App. A, we show how the same result can be recovered using SPA-SALT. Note that in the single-emitter limit where $D_{0}(\mathbf{r})\propto\delta(\mathbf{r}-\mathbf{r}^{\prime})$, Eq. (9) simplifies to $d_{\text{thresh}}=V/Q$, where $V$ is the typical “mode volume” of the resonance, recovering an expression inversely proportional to the Purcell enhancement and the LDOS [LDOS_Laser, LDOS_Laser_2, LDOS_Laser_3]; we comment further on this regime in Sec. 2.3.4.
Once the pumping strength crosses the lasing threshold, the cavity-mode will start to lase with an amplitude $a_{1}$, as depicted in Fig. 1(b). Considering both the zeroth-order and the first-order correction in the small amplitude expansion of Eq. (7), we derive an approximate expression for the lasing intensity
$|a_{1}|^{2}\approx\frac{\Delta d}{d_{\text{thresh}}}\frac{{\int_{\Omega}\varepsilon_{c}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\mathop{}\!\mathrm{d}\Omega}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{2}\,\mathop{}\!\mathrm{d}\Omega}{e_{c}^{-2}\int_{\Omega}D_{0}(\mathbf{r})|\mathbf{e}_{1}(\mathbf{r})|^{4}\,\mathop{}\!\mathrm{d}\Omega}\,$, (10)
where we have assumed that we are operating in the weak pumping limit (just above threshold), where $d\approx d_{\text{thresh}}$ and where $\Delta d=d-d_{\text{thresh}}\ll 1$ represents the pumping strength above threshold. For more details on this derivation, consult App. B. Note that evaluating Eqs. (910) still requires one to solve a linear eigenproblem to determine the modes of the passive cavity, as in SPA-SALT. Nevertheless, assuming a high-$Q$ cavity $(Q\gtrsim 100$), we can eliminate this requirement by exciting the field in the cavity with an external source (e.g., a waveguide coupled to the cavity), since the field in the cavity will be approximately proportional to the passive cavity mode $\mathbf{e}_{1}$ [phot_crys, Rasmus_FOM]. Although this eliminates the requirement of an eigensolve, it does not include information about the efficiency of coupling between the cavity mode and the output channel. In the next section, we will use TCMT to encode the out-coupling into a single, convenient FOM.